X^8=36x^6
Solve for all real solutions

Answer by stanbon(75887)

X^8=36x^6
----
x^8-36x^6 = 0
----
x^6(x^2-36) = 0
x^6(x-6)(x+6) = 0
----
x = 0 or x = 6 or x = -6
-----------
Cheers,
Stan H.
---------------